LINESHAPE ANALYSIS OF CH_{3}F-(ortho-H_{2})_{n} ABSORPTION SPECTRA IN 3000 cm^{−1} REGION IN SOLID para-H_{2}

\underline{Y. MIYAMOTO}, Graduate School of Natural Science and Technology, Okayama University, 3-1-1 Tsushima-naka Okayama 700-8530, Japan; T. MOMOSE, Department of Chemistry, The University of British Columbia, 2036 Main Mall, Vancouver, British Columbia V6T 1Z1, Canada; H. KANAMORI, Department of Physics, Tokyo Institute of Technology, Ohokayama 2-12-1, Tokyo 152-8551, Japan.

FTIR absorption spectra of CH_{3}F-(ortho-H_{2})_{n} clusters in solid para-H_{2} were studied in the C-H stretching region (∼3000 cm^{−1}). As shown previously^{a}, the ν_{3} C-F stretching band at 1040 cm^{−1} of the CH_{3}F-(ortho-H_{2})_{n} (n = 0∼12) gives a distinct regular series of absorption lines in solid para-H_{2}, which correspond to each n-th component of the cluster. On the other hand, broaden single lines were observed in the 3000 cm^{−1} region, which were assigned to the ν_{1} and ν_{4} modes with axial rotation^{b}. We found that their lineshape varied depending on ortho-H_{2} concentration in crystals and could be represented as superposition of the cluster components. Least square fitting based on the CH_{3}F-(ortho-H_{2})_{n} cluster model worked very well to resolve the broaden bands and provided us the linewidth and line shift of resolved lines. The large difference of the linewidth and line shift between the vibrational modes shows the distinct nature of each mode in cold quantum solids.

aK. Yoshioka and D. T. Anderson J. Chem. Phys. \underline{119}, 4731 (2003).

bY. -P. Lee, Y. -J. Wu and J. T. Hougen J. Chem. Phys. \underline{129}, 104502 (2008).